# DeepSieve: Information Sieving via LLM-as-a-Knowledge-Router

### v2

@@ 1 Introduction
However, existing RAG systems still struggle with a fundamental limitation: the lack of an effective information sieving mechanism across the reasoning pipeline. In practice, this deficiency manifests in two critical forms: query-side sieving and source-side sieving. On the query side, most systems treat user queries as atomic units and directly retrieve information without decomposing or analyzing their underlying semantic structure Trivedi et al. (2023); Yao et al. (2023); Lin et al. (2025). This prevents the model from isolating key subgoals or reasoning steps, which is essential for multi-hop or compositional question answering.
@@
To bridge this knowledge gap, we propose DeepSieve, a framework that performs retrieval-augmented reasoning through multi-stage information sieving. As shown in the Figure 1, DeepSieve consists of four key components: question decomposition, thought generation, data-source routing, and recursive reflexion. The pipeline begins by decomposing a complex input query into a set of structured sub-questions using an LLM-based planner Lin et al. (2025); Trivedi et al. (2022). Each sub-question is then independently processed to generate a “thought”, a latent reasoning step that guides retrieval planning Yao et al. (2023); Xu et al. (2023). Based on this thought, DeepSieve identifies the most appropriate source (e.g., API, SQL, RAG corpus) and generates a corresponding action plan. This source-aware routing builds on the idea of using LLMs as controllers for modular tool access Schick et al. (2023); Xu et al. (2025).
@@ 3.1 Experimental Setup
We evaluate on 3 benchmarks: MuSiQue Trivedi et al. (2022), 2WikiMultiHopQA Fu et al. (2021), and HotpotQA Yang et al. (2018), following IRCoT Trivedi et al. (2023) to construct retrieval corpora with both supporting and distractor passages (1,000 dev questions). We use DeepSeek-V3 Guo et al. (2025) and GPT-4o OpenAI et al. (2024) as the backbone LLM. To simulate source heterogeneity, we partition each dataset into local and global segments using LLM-based profiles (see Appendix G). DeepSieve performs subquestion-level routing over these two source modules, allowing modular reasoning across simulated access boundaries. Since existing baselines do not support multi-source retrieval or dynamic source selection, we evaluate them using the original corpus, which is the same as the combination of local and global segments mentioned above, to ensure fair comparison. This setup allows us to isolate the impact of modular routing while maintaining compatibility across methods.
@@
We compare against IRCoT Trivedi et al. (2023), ColBERTv2 Santhanam et al. (2022), HippoRAG Gutierrez et al. (2024), and RAPTOR Sarthi et al. (2024) as representative RAG paradigms baseline. We also include ReAct Yao et al. (2023), ReWOO Xu et al. (2023), Reflexion Madaan et al. (2023), and Chain-of-Thought (CoT) Wei et al. (2022) as reasoning and agentic baselines because DeepSieve utilizes not only RAG algorithm. The details of these baselines and their setups are listed in Appendix F.
@@ 3.2 Main Performance Comparison (RQ1, RQ4)
DeepSeek-V3
Naive RAG (MiniLM) 20.520.520.5 26.126.126.1 26.426.426.4 31.331.331.3 30.430.430.4 42.842.842.8 25.825.825.8 33.433.433.4
ColBERTv2 17.117.117.1 27.227.227.2 32.932.932.9 43.843.843.8 42.942.942.9 57.257.257.2 31.031.031.0 42.742.742.7
HippoRAG 19.419.419.4 27.927.927.9 44.944.944.9 58.958.958.9 40.840.840.8 55.355.355.3 35.035.035.0 47.447.447.4
IRCoT + HippoRAG 21.421.421.4 33.433.433.4 46.546.546.5 63.163.163.1 45.145.145.1 58.958.958.9 37.737.737.7 51.851.851.8
GraphRAG 7.77.77.7 9.79.79.7 30.130.130.1 33.233.233.2 45.945.945.9 55.255.255.2 27.927.927.9 32.732.732.7
RAPTOR 18.218.218.2 28.928.928.9 38.638.638.6 52.152.152.1 52.952.952.9 66.566.566.5 37.537.537.5 50.250.250.2
DeepSieve (Naive RAG) 36.0+14.6↑ 46.8+13.4↑ 62.8+16.3↑ 68.4+5.3↑ 49.049.049.0 61.661.661.6 49.3+11.6↑ 58.9+7.1↑
DeepSieve (GraphRAG) 30.030.030.0+8.6↑ 36.636.636.6+3.2↑ 49.249.249.2+2.7↑ 53.853.853.8 49.449.449.4 61.661.661.6 42.942.942.9+5.2↑ 50.750.750.7
GPT-4o
Naive RAG (MiniLM) 19.419.419.4 27.127.127.1 26.726.726.7 29.829.829.8 29.829.829.8 44.544.544.5 25.325.325.3 33.833.833.8
ColBERTv2 15.515.515.5 26.426.426.4 33.433.433.4 43.343.343.3 43.443.443.4 57.757.757.7 30.830.830.8 42.542.542.5
HippoRAG 19.219.219.2 29.829.829.8 46.646.646.6 59.559.559.5 41.841.841.8 55.055.055.0 35.935.935.9 48.148.148.1
IRCoT + HippoRAG 21.921.921.9 33.333.333.3 47.747.747.7 62.762.762.7 45.745.745.7 56.256.256.2 38.438.438.4 50.750.750.7
GraphRAG 8.78.78.7 10.910.910.9 29.829.829.8 33.333.333.3 44.344.344.3 52.252.252.2 25.625.625.6 31.031.031.0
RAPTOR 18.818.818.8 29.729.729.7 39.339.339.3 51.651.651.6 52.452.452.4 66.366.366.3 36.836.836.8 49.249.249.2
DeepSieve (Naive RAG) 26.7+4.8↑ 36.6+3.3↑ 48.3+0.6↑ 55.255.255.2 49.349.349.3 61.761.761.7 41.4+3.0↑ 51.2+0.5↑
DeepSieve (GraphRAG) 20.320.320.3 31.131.131.1 27.527.527.5 38.938.938.9 44.644.644.6 53.553.553.5 32.732.732.7 43.243.243.2
@@ 3.4 Ablation Study (RQ3)
Setting Decomp. Routing Reflect. HotpotQA 2Wiki. MuSiQue
@@ 4 Related Work
RAG systems augment LLMs with both structured and unstructured knowledge Lewis et al. (2020); Talmor et al. (2021). HippoRAG introduces memory alongside structured retrieval Gutierrez et al. (2024). HippoRAG2 extends it with continual memory using clustering and profiling Gutiérrez et al. (2025). InfuserKI enhances LLMs with knowledge graphs via infusing Wang et al. (2024). AutoSchemaKG automates schema induction for knowledge graph construction from web corpora Bai et al. (2025). These approaches handle heterogeneous memory and structure, but still rely on flat retrieval indexes without per-subquestion navigation.
@@
Agentic systems empower LLMs to reason, plan, and act over multi-step inference chains. Some agentic methods are also used for RAG tasks. ReAct Yao et al. (2023) merges tool use and thought generation in a unified loop. ReWOO Xu et al. (2023) decouples retrieval from reasoning to reduce token cost. MA-RAG Nguyen et al. (2025) introduces agentic collaboration via CoT-based subquerying. AMem Xu et al. (2025) augments LLMs with lifelong memory and dynamic memory selection during inference. In contrast, DeepSieve builds upon this line of work by introducing an explicit modular planner-router-retriever loop, enabling fine-grained reasoning over decomposed subquestions, recursive reflexion over failures, and heterogeneous tool dispatch.
@@ 5 Conclusion
Limitations and Future Work
While DeepSieve demonstrates strong performance in modular reasoning and heterogeneous source routing, there remain two key directions for further enhancement: First, the current routing mechanism selects only a coarse-grained (tool, source) pair for each subquestion. This limits the system’s ability to leverage fine-grained configurations, such as tool-specific parameters (e.g., retrieval depth, temperature, API mode) or function-level APIs. Future work could extend the action space to support parameterized tool selection Xu et al. (2023), enabling more adaptive behavior, utilizing cache to save sources Zhu et al. (2025), and cost-aware decisions at inference time Mei et al. (2025a). Second, although DeepSieve supports modular integration of private sources, it treats all subquestions uniformly across users. In real-world settings, however, different users may have personalized knowledge graphs Du et al. (2025), access patterns, or preferences. A promising direction is to incorporate personalized routing and memory modules, allowing the LLM to learn user-specific retrieval paths, preferred sources, or task priors, thus enabling long-term adaptation and user-centric QA behavior. We believe these extensions will further enhance the controllability and personalization of RAG systems in both industrial and research applications.
@@ Appendix E Pseudocode Implementation
[…]
@@
while $j<C.max\_reflexion\_attempts$ and not $is\_success$ do
     if $C.use\_routing$ then
@@
$a_{i},is\_success\leftarrow\text{ExtractAnswer}(a_{candidate})$ // Stage III: Observe if not $is\_success$ and $C.use\_reflexion$ then
@@
[…]
@@ Results.
Table 5: Deepsieve vs. Expert Oracle performance (Exact Match ||| F1). The underline means that the results are calculated in an ideal setting which is not feasible in real scene. Setting SQL RAG Overall Expert Oracle (S1) 45.0 ||| 46.2 32.4 ||| 41.8 38.7 ||| 44.0 DeepSieve (S2) 50.8 ||| 52.0 48.3 ||| 59.1 35.6 ||| 40.1

### v3

@@ 1 Introduction
However, existing RAG systems still struggle with a fundamental limitation: the lack of an effective information sieving mechanism across the reasoning pipeline. In practice, this deficiency manifests in two critical forms: query-side sieving and source-side sieving. On the query side, most systems treat user queries as atomic units and directly retrieve information without decomposing or analyzing their underlying semantic structure Trivedi et al. (2023); Yao et al. (2023); Lin et al. (2025b). This prevents the model from isolating key subgoals or reasoning steps, which is essential for multi-hop or compositional question answering.
@@
To bridge this knowledge gap, we propose DeepSieve, a framework that performs retrieval-augmented reasoning through multi-stage information sieving. As shown in the Figure 1, DeepSieve consists of four key components: question decomposition, thought generation, data-source routing, and recursive reflexion. The pipeline begins by decomposing a complex input query into a set of structured sub-questions using an LLM-based planner Lin et al. (2025b); Trivedi et al. (2022). Each sub-question is then independently processed to generate a “thought”, a latent reasoning step that guides retrieval planning Yao et al. (2023); Xu et al. (2023b). Based on this thought, DeepSieve identifies the most appropriate source (e.g., API, SQL, RAG corpus) and generates a corresponding action plan. This source-aware routing builds on the idea of using LLMs as controllers for modular tool access Schick et al. (2023); Xu et al. (2025).
@@ 3.1 Experimental Setup
We evaluate on 3 benchmarks: MuSiQue Trivedi et al. (2022), 2WikiMultiHopQA Ho et al. (2020), and HotpotQA Yang et al. (2018), following IRCoT Trivedi et al. (2023) to construct retrieval corpora with both supporting and distractor passages (1,000 dev questions). We use DeepSeek-V3 Guo et al. (2025) and GPT-4o OpenAI et al. (2024) as the backbone LLM. To simulate source heterogeneity, we partition each dataset into local and global segments using LLM-based profiles (see Appendix G). DeepSieve performs subquestion-level routing over these two source modules, allowing modular reasoning across simulated access boundaries. Since existing baselines do not support multi-source retrieval or dynamic source selection, we evaluate them using the original corpus, which is the same as the combination of local and global segments mentioned above, to ensure fair comparison. This setup allows us to isolate the impact of modular routing while maintaining compatibility across methods.
@@
We compare against IRCoT Trivedi et al. (2023), ColBERTv2 Santhanam et al. (2022), HippoRAG Gutierrez et al. (2024), and RAPTOR Sarthi et al. (2024b) as representative RAG paradigms baseline. We also include ReAct Yao et al. (2023), ReWOO Xu et al. (2023a), Reflexion Shinn et al. (2023), and Chain-of-Thought (CoT) Wei et al. (2022) as reasoning and agentic baselines because DeepSieve utilizes not only RAG algorithm. The details of these baselines and their setups are listed in Appendix F.
@@ 3.2 Main Performance Comparison (RQ1, RQ4)
DeepSeek-V3 Naive RAG (MiniLM) $20.5$ $26.1$ $26.4$ $31.3$ $30.4$ $42.8$ $25.8$ $33.4$
ColBERTv2 $17.1$ $27.2$ $32.9$ $43.8$ $42.9$ $57.2$ $31.0$ $42.7$
HippoRAG $19.4$ $27.9$ $44.9$ $58.9$ $40.8$ $55.3$ $35.0$ $47.4$
IRCoT + HippoRAG $21.4$ $33.4$ $46.5$ $63.1$ $45.1$ $58.9$ $37.7$ $51.8$
GraphRAG $7.7$ $9.7$ $30.1$ $33.2$ $45.9$ $55.2$ $27.9$ $32.7$
RAPTOR $18.2$ $28.9$ $38.6$ $52.1$ $52.9$ $66.5$ $37.5$ $50.2$
DeepSieve (Naive RAG) 36.0+14.6↑ 46.8+13.4↑ 62.8+16.3↑ 68.4+5.3↑ $49.0$ $61.6$ 49.3+11.6↑ 58.9+7.1↑
DeepSieve (GraphRAG) 30.030.0+8.6↑ 36.636.6+3.2↑ 49.249.2+2.7↑ $53.8$ $49.4$ $61.6$ 42.942.9+5.2↑ $50.7$
GPT-4o Naive RAG (MiniLM) $19.4$ $27.1$ $26.7$ $29.8$ $29.8$ $44.5$ $25.3$ $33.8$
ColBERTv2 $15.5$ $26.4$ $33.4$ $43.3$ $43.4$ $57.7$ $30.8$ $42.5$
HippoRAG $19.2$ $29.8$ $46.6$ $59.5$ $41.8$ $55.0$ $35.9$ $48.1$
IRCoT + HippoRAG $21.9$ $33.3$ $47.7$ $62.7$ $45.7$ $56.2$ $38.4$ $50.7$
GraphRAG $8.7$ $10.9$ $29.8$ $33.3$ $44.3$ $52.2$ $25.6$ $31.0$
RAPTOR $18.8$ $29.7$ $39.3$ $51.6$ $52.4$ $66.3$ $36.8$ $49.2$
DeepSieve (Naive RAG) 26.7+4.8↑ 36.6+3.3↑ 48.3+0.6↑ $55.2$ $49.3$ $61.7$ 41.4+3.0↑ 51.2+0.5↑
DeepSieve (GraphRAG) $20.3$ $31.1$ $27.5$ $38.9$ $44.6$ $53.5$ $32.7$ $43.2$
@@ 3.4 Ablation Study (RQ3)
Setting Decomp. Routing Reflexion HotpotQA 2Wiki. MuSiQue
@@ 4 Related Work
RAG systems augment LLMs with both structured and unstructured knowledge Lewis et al. (2020); Talmor et al. (2021). HippoRAG introduces memory alongside structured retrieval Gutierrez et al. (2024). HippoRAG2 extends it with continual memory using clustering and profiling Gutiérrez et al. (2025). InfuserKI enhances LLMs with knowledge graphs via infusing Wang et al. (2024). AutoSchemaKG automates schema induction for knowledge graph construction from web corpora Bai et al. (2025). These approaches handle heterogeneous memory and structure, but still rely on flat retrieval indexes without per-subquestion navigation. Beyond retrieval, routing-based agent frameworks can explicitly decide which knowledge/tool/agent to invoke at each step; for example, Sun propose fast–slow routing for visual agents to switch between lightweight and deliberative reasoning modules Sun et al. (2025).
@@
Agentic systems empower LLMs to reason, plan, and act over multi-step inference chains. Some agentic methods are also used for RAG tasks. ReAct Yao et al. (2023) merges tool use and thought generation in a unified loop. ReWOO Xu et al. (2023b) decouples retrieval from reasoning to reduce token cost. MA-RAG Nguyen et al. (2025) introduces agentic collaboration via CoT-based subquerying. Recent works have also explored optimizing agentic RAG systems through caching mechanisms Lin et al. (2025a) to reduce latency and cost. In contrast, DeepSieve builds upon this line of work by introducing an explicit modular planner-router-retriever loop, enabling fine-grained reasoning over decomposed subquestions, recursive reflexion over failures, and heterogeneous tool dispatch.
@@ 5 Conclusion
Limitations
While DeepSieve demonstrates strong performance in modular reasoning and heterogeneous source routing, there remain directions for further enhancement: First, the current routing mechanism selects only a coarse-grained (tool, source) pair for each subquestion. This limits the system’s ability to leverage fine-grained configurations, such as tool-specific parameters (e.g., retrieval depth, temperature, API mode) or function-level APIs. Future work could extend the action space to support parameterized tool selection, enabling more adaptive behavior and cost-aware decisions at inference time (Mei et al., 2025c).
Second, although DeepSieve supports modular integration of private sources, it treats all subquestions uniformly across users. In real-world settings, however, different users may have personalized knowledge graphs, access patterns, or preferences. A promising direction is to incorporate personalized routing and memory modules, allowing the LLM to learn user-specific retrieval paths, preferred sources, or task priors, thus enabling long-term adaptation and user-centric QA behavior.
Third, the interpretability of the agentic decision-making process is critical for user trust and system debugging. While DeepSieve effectively routes queries, the rationale behind specific routing choices can be opaque. Future work could incorporate agentic explainer frameworks, such as SAGE (Han et al., 2025), to provide granular interpretations of the router’s internal features and verify the logic behind its source selection and decomposition steps.
Furthermore, to further optimize routing precision, we plan to move beyond surface-level confidence scores. Recent findings on concept depth (Jin et al., 2025) reveal how LLMs acquire knowledge across layers. Future iterations of DeepSieve could leverage these layer-wise signals to distinguish between well-mastered internal knowledge and hallucinations, triggering retrieval tools only when the model’s internal concept depth is insufficient. We believe these extensions will further enhance the controllability and personalization of RAG systems in both industrial and research applications.
Finally, our current evaluation primarily focuses on standard multi-hop question answering benchmarks. We have not yet explored the applicability of DeepSieve in more dynamic, interactive, or simulation-based environments. Future work could investigate how our routing mechanism adapts to complex scenarios such as computer-use agent (Mei et al., 2025a), game prototyping (Li et al., 2025b, a, 2026) ,digital twin persona simulation (Du et al., 2025, ) and activity recognition (Li et al., 2020), where information retrieval must align with evolving states or specific character personas.
Acknowledgements
We would like to thank Mingyu Jin for his contributions to this work.
@@ Appendix E Pseudocode Implementation
Algorithm: The DeepSieve Pipeline Input: Query $Q$, Set of source modules $S$, Config $C$ Output: Final answer $\hat{A}$ $M\leftarrow\text{InitializeMemory()}$ // memory for results & failures if $C.decompose$ then
@@
while j<C.max_reflexion_attemptsj<C.max\_reflexion\_attempts and not is_successis\_success do
     if C.use_routingC.use\_routing then
@@
$a_{i},is\_success\leftarrow\text{ExtractAnswer}(a_{candidate})$ // Stage III: Observe if not is_successis\_success and C.use_reflexionC.use\_reflexion then
@@
return $\hat{A}$ This design is a significant advantage as it was crucial for conducting the ablation studies presented in our experiments. The algorithm’s versatility is highlighted by the if C.use_routing block, which shows how the framework can execute its primary logic for heterogeneous sources or gracefully degrade to a standard "RAG-only" baseline operating on a single merged source. Algorithm: Helper Functions for DeepSieve Function Decompose($Q$): $prompt\leftarrow\text{CreateDecompositionPrompt}(Q)$ $response\leftarrow\text{LLM}(prompt)$ return $\text{ParseSubqueryDAG}(response)$ Function Route($q$, $S$, FailsFails): $prompt\leftarrow\text{CreateRoutingPrompt}(q,S,Fails)$ $response\leftarrow\text{LLM}(prompt)$ return $\text{SelectSourceFrom}(response,S)$ Function ExtractAnswer($a_{candidate}$): $prompt\leftarrow\text{CreateExtractionPrompt}(a_{candidate})$ $response\leftarrow\text{LLM}(prompt)$ $(a,s)\leftarrow\text{ParseAnswerAndSuccess}(response)$ return $(a,s)$ Function Fuse($Q$, $M$): $prompt\leftarrow\text{CreateFusionPrompt}(Q,M.\text{successes})$ $response\leftarrow\text{LLM}(prompt)$ return $\text{ParseFinalAnswer}(response)$ The reasoning process begins with an explicit planning step, where the Decompose function generates a structured Directed Acyclic Graph (DAG) of subqueries. A key strength of this approach is that it makes the reasoning plan transparent, machine-readable, and capable of modeling the complex, non-linear dependencies required to solve multi-hop questions. The execution of this plan is both dynamic and adaptive. For each subquery, the Route function intelligently selects a knowledge source, while the while loop embodies the Reflexion stage, providing a robust self-correction mechanism. If a retrieval is unsuccessful, the failure is logged to memory, allowing the system to attempt a different route on the next iteration. This entire workflow is powered by the helper functions in the helper function algorithm, which abstract the core cognitive tasks into distinct, prompt-driven LLM calls. By encapsulating operations like Decompose and Fuse into modular functions, the framework adheres to the "LLM-as-a-Controller" paradigm. This has the advantage of making the system’s logic transparent and easy to modify, as the behavior of each stage can be fine-tuned simply by refining its corresponding prompt without altering the core control flow code.
@@ Results.
Table 5: DeepSieve vs. Expert Oracle performance (Exact Match $|$ F1). The underline means that the results are calculated in an ideal setting which is not feasible in real scene. Setting SQL RAG Overall Expert Oracle (S1) 45.0 $|$ 46.2 32.4 $|$ 41.8 38.7 $|$ 44.0 DeepSieve (S2) 50.8 $|$ 52.0 48.3 $|$ 59.1 35.6 $|$ 40.1
@@
Appendix J Real-World Heterogeneous Experiment: MedQA + CaseHOLD
Motivation.
While the experiments in former sections validate routing under controlled conditions, a key question still remains: can DeepSieve handle truly heterogeneous, real-world domains that are naturally incompatible? To address this, and to verify that our gains are not artifacts of artificial partitioning, we conducted an additional evaluation combining two distinct domain-specific corpora: MedQA (Medical) Jin et al. (2021) and CaseHOLD (Legal) Zheng et al. (2021).
Setup.
We constructed a unified heterogeneous benchmark by mixing queries and corpora from MedQA and CaseHOLD. This setup mimics a realistic multi-tenant or multi-domain system where a single entry point must correctly route specialized queries to their respective knowledge bases (Medical vs. Legal) without cross-domain pollution.
Results.
Table 6 presents the performance of DeepSieve against standard and Graph RAG baselines. DeepSieve significantly outperforms the baselines, achieving a +50% improvement in Exact Match (EM) over Naive RAG. Method EM F1 Naive RAG 16.00 23.45 DeepSieve (Naive RAG) 24.00 34.39 Graph RAG 18.00 27.14 DeepSieve (Graph RAG) 22.00 32.82 Table 6: Performance comparison on the combined MedQA (Medical) and CaseHOLD (Legal) setting. DeepSieve demonstrates superior adaptability to truly heterogeneous sources.
Analysis.
The substantial performance gap highlights the critical role of source-aware routing in real-world scenarios. Standard RAG systems often suffer from "context pollution," retrieving irrelevant legal precedents for medical queries (or vice versa) due to vector space collisions. DeepSieve effectively isolates these domains via its routing stage.
Latency Note.
We also evaluated the wall-clock latency in this setting. DeepSieve averages 0.058s per query, compared to 0.026s for Naive RAG. The slight overhead ($\sim$0.03s) incurred by the decomposition and routing steps is negligible for real-time applications, especially considering the significant gains in accuracy and the reduction in generated tokens.